OpenStudy (ashley7777):

sin^2x-sinx-1=0 solve for x

OpenStudy (freckles):

this is a quadratic in terms of sin(x) replace sin(x) with u it might make it easier for you see that

OpenStudy (freckles):

\[(\sin(x))^2-\sin(x)-1=0 \\ u^2-u-1=0\] solve for u then replace u with sin(x) and solve for x